OpenStudy (anonymous):

3x-y=7 2x-2y=2 Find for x and y?

OpenStudy (arabpride):

Alright. *^.^* Do you know how to simplify?

OpenStudy (anonymous):

No/:

OpenStudy (arabpride):

It's ok. Would you like me to show you the steps?

OpenStudy (anonymous):

@WonderlandKy , this is the same process as the other one I showed you. Solve the first equation for either x or y (your choice), and THEN substitute/plug that value you just got for x or y into either of the two equations to solve for the other variable.

OpenStudy (anonymous):

Yes please c: I had asked a question before and had been going off that example but it closed and won't let me open it again/:

OpenStudy (anonymous):

@TOMWALK yes, I had gone off that example for a few others but it wouldn't let me open the question back up/: Thank you so much!c:

OpenStudy (anonymous):

Here's the previous example: 13x-2y=70 2x-2y=4 1) 13x-2y=70 --> rearrange to get 2y = 13x-70 Substitute that value for 2y into the 2nd equation: 2x-2y=4 2x - (13x-70) = 4 2x - 13x + 70 = 4 -11x = -66 x = 6 2) Now substitute this value for x into EITHER of the two original equations to solve for y... 13x-2y = 70 13(6) - 2y = 70 78 - 2y = 70 -2y = -8 y = 4 Solution set: {6,4}

OpenStudy (anonymous):

@TOMWALK Thank you!:D

OpenStudy (anonymous):

Have you studied solving systems of equations by other methods other than substitution, such as elimination? There are other ways to solve these types of systems of equations. I could show you.

OpenStudy (arabpride):

solving for y: 3x-y=7 -3x -3x -y = -3x + 7 /-1 /-1 ~~~~~~~~~~ y = 3x - 7 ~~~~~~~~~~ now to solve for x on the same original equation: 3x-y=7 +y +y 3x = y + 7 /3 /3 ~~~~~~~~~~~~ x = y/3 + 7/3 ~~~~~~~~~~~~ *********************************************************************** solving for y: 2x-2y=2 -2x -2x -2y = -2x + 2 /-2 /-2 ~~~~~~~~~ y = x - 1 ~~~~~~~~~ now solving for x: 2x-2y=2 +2y +2y 2x = 2y + 2 /2 /2 ~~~~~~~~ x = y + 1 ~~~~~~~~

OpenStudy (anonymous):

Yes, please c:

OpenStudy (anonymous):

Thank you @arabpride so much c:

OpenStudy (arabpride):

You're Very Welcome *^.^*

OpenStudy (anonymous):

You can also solve these types of problems by simply adding them to eliminate. For example: You start with: 3x-y=7 2x-2y=2 MULTIPLYING THE ENTIRE TOP EQUATION BY -2 WILL RESULT IN THE Y-TERM CANCELLING OUT WHEN YOU ADD THE TWO EQUATIONS TOGETHER: -2(3x-y=7) +2x-2y=2 -6x+2y=-14 +2x-2y=2 -4x=-12 x = 3 NOW, take this X value and plug it/substitute it into EITHER of the original two equations to solve for Y.

OpenStudy (anonymous):

Remember that you actually need to come up with an integer solution set. So X = some integer and Y = some integer. Your final answer needs to be in that format. In other words, your final answer for X should not contain the variable Y, and vice versa. It should just be a number (an integer).